OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!! dora wrote the equation of the circle in the center radius form as (x - p)² + (y - q)² = 16. what is the value of p+q?

OpenStudy (anonymous):

But if you only have the radius, the center point could be anywhere on the point o-o

OpenStudy (anonymous):

does this help?

OpenStudy (anonymous):

verymuch.

OpenStudy (anonymous):

the answer is 6 because 3+3=6. Ask me if you have questions.

OpenStudy (anonymous):

I do umm how did you get 3+3=6?

OpenStudy (anonymous):

It's so obvious ;)

OpenStudy (anonymous):

im still confused

OpenStudy (anonymous):

well, if you look at the graph, the center point is at (3,3)

OpenStudy (anonymous):

yeah but shouldn't it match to 16?

OpenStudy (anonymous):

p and q are the coordinates for the center in x and y

OpenStudy (anonymous):

16 is just the indication of r^2 which means that the radius is 4

OpenStudy (anonymous):

Ok

OpenStudy (anonymous):

but you already have that in the graph

OpenStudy (anonymous):

thats what they gave me

OpenStudy (anonymous):

From the center, draw the 2 lines perpendicularly to x axis, y axis! It called coordinate ( y, p)

OpenStudy (anonymous):

=> y + p = ...

OpenStudy (anonymous):

You're no fun :P

OpenStudy (anonymous):

so whats the answer here...